Consider the following two scenarios. The taxpayer sells the property acquired in April 2019 for $800,000 (rental real estate, 27.5 year depreciation) on December 31, 2019 or December 31, 2020 for $1,000,000. Assume the taxpayer is in the 37 percent tax bracket.

What is the gain on the sale? Show Calculations.
Answer #1

Consider the following two scenarios. The taxpayer sells the property acquired in April 2019 for $800,000 (rental real estate, 27.5 year depreciation) on December 31, 2019 or December 31, 2020 for $1,000,000. Assume the taxpayer is in the 37 percent tax bracke

Gain on the sale =$1000000-800000(27.5/12/100)=181666.7-37% Tax=176822.2